OpenStudy (anonymous):

Is the sequence {10,6,2,-2 -6...... } arithmetic? If so, what is the common difference?

OpenStudy (anonymous):

c.d=a2-a1=6-10=-4

OpenStudy (anonymous):

how is that ?

OpenStudy (anonymous):

c.d in an A.P=an-an-1 c.d=any term- preceding term c.d=6-10=2-6=-2-2=-4

OpenStudy (anonymous):

okay gotchu

OpenStudy (anonymous):

yw